### v1

# Multimodal atmospheric super-resolution with deep generative models

Abstract
Score-based diffusion modeling is a generative machine learning algorithm that can be used to sample from complex distributions. They achieve this by learning a score function, i.e., the gradient of the log-probability density of the data, and reversing a noising process using the same. Once trained, score-based diffusion models not only generate new samples but also enable zero-shot conditioning of the generated samples on observed data. This promises a novel paradigm for data and model fusion, wherein the implicitly learned distributions of pretrained score-based diffusion models can be updated given the availability of online data in a Bayesian formulation. In this article, we apply such a concept to the super-resolution of a high-dimensional dynamical system, given the real-time availability of low-resolution and experimentally observed sparse sensor measurements from multimodal data. Additional analysis on how score-based sampling can be used for uncertainty estimates is also provided. Our experiments are performed for a super-resolution task that generates the ERA5 atmospheric dataset given sparse observations from a coarse-grained representation of the same and/or from unstructured experimental observations of the IGRA radiosonde dataset. We demonstrate accurate recovery of the high dimensional state given multiple sources of low-fidelity measurements. We also discover that the generative model can balance the influence of multiple dataset modalities during spatiotemporal reconstructions.
Keywords Generative machine learning  $\cdot$ Multimodal data fusion  $\cdot$ Atmospheric super-resolution
1 Introduction
Several real-world forecasting applications involve high-dimensional dynamical systems that exhibit multiscale behavior. A classical example of such a system is the Earth’s atmosphere which exhibits rich spatiotemporal behavior [1, 2, 3]. The modeling and forecasting of such a system is complicated by the challenges involved in simulating and observing them. In the case of the atmosphere, a variety of instrumentation are leveraged to partially sense the atmosphere [4, 5]. These partial observations are used to enhance understanding of atmospheric processes and improve forecasts, for example, by correcting dynamical models with parameterizations or improved initial conditions[6, 7, 7, 8, 9]. However, when spatial and temporal resolution are limited for sensing, an inverse problem emerges for recovering the full state of the dynamical system in the presence of partial observations. The dynamical models themselves, based on numerical discretizations of nonlinear partial differential equations, require vast computational resources for generating forecasts [10]. Ultimately, the most accurate forecasting paradigm relies on data and model fusion, wherein first-principles-based numerical models are adaptively corrected in real-time within a Bayesian statistics formulation [11, 9]. In such an approach, the numerical models are assumed to provide a prior (or background) state which is updated given partial observations that are used to compute a likelihood. In this article, we investigate the aforementioned computational framework from the perspective of generative machine learning. Generative machine learning [12], in contrast with discriminative models, represents a paradigm where data is used to learn an underlying probability density function from which novel samples can be drawn (or ‘generated’) [13, 14, 15]. In particular, we continue to frame the concept of data and model fusion within a Bayesian setting but leverage deep learning for implicit learning and sampling from an underlying distribution that can be updated in real-time given sparse observations.
Given a dataset sampled from an unknown distribution, the goal of a generative model is to approximate this distribution with a model that allows sampling, inference, or density estimation [12]. Broadly, generative models fall into two categories: explicit density models, which define a likelihood function and aim to estimate it directly, and implicit density models, which focus on generating samples without requiring an explicit form of the likelihood. A state-of-the-art generative machine learning algorithm is the so-called score-based diffusion model [16, 17, 18, 19] – a focus of this study. In this approach, one starts with data (such as images) and gradually adds noise over time through a forward stochastic process—typically a diffusion process—until the data becomes nearly indistinguishable from pure Gaussian noise. This forward process is mathematically described using stochastic differential equations (SDEs), and it defines a continuous trajectory from clean data to random noise [20]. It is important to note that score-based models learn to approximate, using a neural network, the score function—the gradient of the log-probability density of the data at various noise levels. This score function tells the model how to adjust a noisy sample to make it more likely under the data distribution. Finally, to generate new data, the model starts with a sample of pure noise and simulates the reverse-time SDE using the learned score function, effectively “denoising" the sample step-by-step. In essence, a trained score-based diffusion model implicitly learns an unknown distribution given collected samples. Furthermore, score-based models can also be used to adapt samples given sparse observations of the generated samples. This is achieved using ‘zero-shot’ score matching [21, 22], where a pretrained score-based diffusion model is conditioned on new, sparse observations—without retraining the model. The core idea is to modify the sampling process using observations (constraints) to guide the generation, effectively sampling from a conditional distribution.
In this article, we develop score-based diffusion models for generating realistic states of a high-dimensional multiscale system (the atmosphere) from sampled noise. Subsequently, we utilize our pretrained model for rapidly generating samples that assimilate observations from coarse grids or from different sensor measurements. We outline specific contributions in the following
  1. 1.
    We train a score-based diffusion model using ERA5 reanalysis data [23] to generate realistic samples corresponding to the global atmosphere for 69 variables on a 1.4-degree resolution grid.
  2. 2.
    We perform zero-shot score-matching using sparse observations to demonstrate a super-resolution task for full-state recovery with quantified uncertainty.
  3. 3.
    We perform zero-shot score-matching using partial observations of different datasets to demonstrate a multifidelity super-resolution task with quantified uncertainty.
Our results demonstrate that score-based generative models can be used to generate high-dimensional states of dynamical systems (i.e., in a super-resolution task) from a variety of data sources without retraining using zero-shot score-matching. Moreover, we also demonstrate efficient ‘data-fusion’ where zero-shot samples can balance the influence of multiple sources of data that are partially observed. This allows for the utilization of different sources of data available at different spatial resolutions for posterior updates during the generative sampling process.
1.1 Related work
In recent studies, deep learning-based models have provided exciting results in the state-reconstruction of high-dimensional multiscale dynamical systems [24, 25, 26]. For example, in applications related to the reconstruction of turbulent flows and for magnetic resonance imaging data, customized physics-informed deep learning methods have led to the remarkable recovery of fine-scaled features from coarse-grained observations [27, 28]. Most work in this emergent area of data-driven super-resolution has relied on structured grid representations of both the coarse and fine representations of the state which has allowed for the use of convolutional neural network techniques [29, 30]. For unstructured representations, graph neural network applications have also been used with significant success [31]. Methods that bridge convolutional architectures and arbitrary locations of sparse observations using Voronoi tessellation have also led to impressive state reconstructions [32]. However, a significant majority of these super-resolution demonstrations possess two limitations: first, they frequently rely on coarse-grid representations extracted from the fine state (i.e., they are obtained from the same dataset). Second, the proposed models need retraining when the source or representation of coarse-grid data is significantly changed for example when it is obtained from a different dataset. Finally, since super-resolution is an inverse problem, several deterministic techniques face limitations in deployments without some form of uncertainty quantification.
Some deep learning super-resolution methods that deploy Uncertainty Quantification (UQ) have leveraged probabilistic neural networks for quantifying aleatoric uncertainty [33] and deep ensembles for quantifying epistemic uncertainty [34]. However, the former typically relies on the assumption of a Gaussian likelihood and the latter requires significant computational costs for the parallel training of multiple neural networks. Stochastic weight averaging [35] is a competitive alternative for computationally efficient UQ in arbitrary deep learning tasks but requires restrictive assumptions on the use of the optimizer as well as the nature of the loss surface (requiring an approximately quadratic loss surface upon convergence) [36]. This motivates the use of generative modeling for our present application.
The landscape of generative modeling for complex physical systems has been reshaped by the advent of deep learning, with score-based diffusion models emerging as a particularly potent paradigm. This evolution began with the foundational work on diffusion probabilistic models [37], which was significantly advanced by Denoising Diffusion Probabilistic Model (DDPM) [38], demonstrating its ability for high-fidelity image synthesis. Currently, the development of score-based generative models, which learn the log density gradient of the data distribution, provided a powerful alternative [39]. These two perspectives were unified and generalized through the framework of stochastic differential equations (SDEs) [40], establishing a continuous-time formulation for the diffusion and reverse-time generation processes. Further refinements to the design space, including preconditioning, noise scheduling, and sampling, led to Elucidating Diffusion Models (EDM), which achieved state-of-the-art results with improved efficiency [41]. More recently, Flow Matching has been introduced as a simulation-free training paradigm for continuous normalizing flows that is more stable and efficient, generalizing diffusion paths and allowing for novel transport paths, such as those based on optimal transport [42, 43, 44, 45]. These powerful generative priors have been adeptly applied to solve a wide range of inverse problems, often in a zero-shot or plug-and-play manner. Techniques such as Diffusion Posterior Sampling (DPS) [46] and Denoising Diffusion Restoration Models (DDRM) [47] enable sampling from the posterior distribution conditioned on measurements, facilitating tasks like super-resolution, inpainting, and deblurring without retraining the base model [48, 49, 50].
In this work, we utilized generative deep learning using EDM models to enable the super-resolution of high-dimensional states from dynamical systems from arbitrary coarse-grid observations without retraining. We also jointly provide UQ without the aforementioned restrictive assumptions through sampling from the generative process. Before proceeding, we review related work that has leveraged generative models for super-resolution of high-dimensional dynamical systems.
In the domain of atmospheric and climate science, these generative models are revolutionizing weather forecasting and data assimilation. Numerous studies have demonstrated their use for downscaling and super-resolution [51, 52, 53, 54, 55], probabilistic and ensemble forecasting [56, 57, 58, 59, 60], and data assimilation [61, 62, 63]. Specifically, models like SEEDS [56] have shown the ability to generate large, skillful ensembles for weather forecasting at a fraction of the computational cost of traditional methods. [62] demonstrates the viability of score-based data assimilation for incorporating sparse weather station data at kilometer scales, a task highly relevant to our work. Other similar applications include precipitation nowcasting [64, 65], emulation of Earth system models (DiffESM - [66]), and generating climate simulations [67]. These notable climate diffusion models along with FuXi-Extreme [68], DYffusion [69] and Appa [70] underscore the transformative potential of generative models to handle the high dimensionality, uncertainty, and multiscale nature of atmospheric dynamics. In this study, we utilize diffusion models for downscaling from sparse observational data, while investigating the influence of multiple sources of data at different fidelities.
2 Methods
Diffusion models represent a new paradigm in generative modeling, achieving remarkable success in generating high-fidelity data across various modalities, including images, audio, and more [38, 40, 71]. These models are inspired by concepts from non-equilibrium thermodynamics [37] and typically involve a two-stage process: a forward process that gradually injects noise into data, and a learned reverse process that aims to denoise and generate data. We develop the mathematical formulation of the various steps of the algorithm in the following.
2.1 Forward and Reverse Diffusion Processes
The forward diffusion process systematically corrupts an initial data sample $\mathbf{x}_{0}$ (drawn from the true data distribution $q(\mathbf{x})$) with noise. In the continuous-time formulation central to our work, this can be described by a forward stochastic differential equation (SDE) [41, 40]:
$d\mathbf{x}=\mathbf{f}(\mathbf{x},t)dt+g(t)d\mathbf{w}$ (1)
The drift term $\mathbf{f}(\mathbf{x},t)$ governs the deterministic evolution of the data, while the diffusion term $g(t)$ scales the noise added via a Wiener process $d\mathbf{w}$. If the drift is zero, often referred to as the Variance Preserving (VP) SDE, the process simplifies to the addition of Gaussian noise with a specific standard deviation $\sigma$:
$\mathbf{x}(\sigma)=\mathbf{x}_{0}+\sigma\mathbf{n}$ (2)
where $\mathbf{n}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$. The noise level $\sigma$ ranges from $\sigma_{\text{min}}\approx 0$ (no noise) to $\sigma_{\text{max}}$ (pure noise), at which point the data distribution $p(\mathbf{x}(\sigma_{\text{max}}))$ converges to a simple isotropic Gaussian prior. The generative power of these models comes from learning to reverse this process which has a corresponding reverse-time SDE [72]. Solving this reverse SDE requires the score function, $\nabla_{\mathbf{x}(\sigma)}\log p(\mathbf{x}(\sigma))$, which points in the direction of increasing data density. For the Gaussian noise model, a key relationship is:
$\nabla_{\mathbf{x}(\sigma)}\log p(\mathbf{x}(\sigma))=-\frac{\mathbb{E}_{\mathbf{x}_{0}|\mathbf{x}(\sigma)}[\mathbf{x}(\sigma)-\mathbf{x}_{0}]}{\sigma^{2}}$ (3)
A neural network, $D_{\theta}$, is trained to approximate this score. If the network is trained to predict the clean data, $D_{\theta}(\mathbf{x}(\sigma);\sigma)\approx\mathbf{x}_{0}$, its output can be used to estimate the score:
$\nabla_{\mathbf{x}(\sigma)}\log p(\mathbf{x}(\sigma);\sigma)\approx\frac{D_{\theta}(\mathbf{x}(\sigma);\sigma)-\mathbf{x}(\sigma)}{\sigma^{2}}$ (4)
Alternatively, if the network predicts the noise, $\boldsymbol{\epsilon}_{\theta}(\mathbf{x}(\sigma);\sigma)\approx\mathbf{n}$:
$\nabla_{\mathbf{x}(\sigma)}\log p(\mathbf{x}(\sigma);\sigma)\approx-\frac{\boldsymbol{\epsilon}_{\theta}(\mathbf{x}(\sigma);\sigma)}{\sigma}$ (5)
This allows the reverse process to be formulated as an ordinary differential equation (ODE), known as the probability flow ODE, which can be solved numerically to generate samples.
2.2 EDM Formulation and Training
The Elucidating Diffusion Model (EDM) framework [41] provides a principled design space for diffusion models. A key aspect is network preconditioning, where the overall denoising model $F_{\theta}$ is defined as:
$F_{\theta}(\mathbf{x}(\sigma);\sigma)=c_{\text{skip}}(\sigma)\mathbf{x}(\sigma)+c_{\text{out}}(\sigma)D_{\theta}\left(\frac{\mathbf{x}(\sigma)}{c_{\text{in}}(\sigma)};c_{\text{noise}}(\sigma)\right)$ (6)
where $D_{\theta}$ is the neural network parameterized by $\theta$. The scaling factors ($c_{\text{skip}},c_{\text{out}},c_{\text{in}},c_{\text{noise}}$) are crucial for ensuring the network operates on inputs with approximately unit variance. We use the specific coefficient definitions from Table 1 in [41]. The training objective is a weighted mean squared error:
$L(\theta)=\mathbb{E}_{\mathbf{x}_{0},\mathbf{n},\sigma}\left[\lambda(\sigma)||\mathbf{x}_{0}-F_{\theta}(\mathbf{x}_{0}+\sigma\mathbf{n};\sigma)||^{2}\right]$ (7)
where $\sigma$ is sampled from a distribution $p(\sigma)$ and $\lambda(\sigma)$ is a weighting function, typically $\lambda(\sigma)=1/c_{\text{out}}(\sigma)^{2}$, that balances the loss contribution across different noise levels.
2.3 Sampling
Once trained, the model $F_{\theta}$ generates samples by numerically solving the reverse ODE. Starting from a noise sample $\mathbf{x}_{\sigma_{\text{max}}}\sim\mathcal{N}(\mathbf{0},\sigma_{\text{max}}^{2}\mathbf{I})$, an ODE solver iteratively refines the sample across a sequence of decreasing noise levels $\sigma_{i}$. A first-order Euler step from $\mathbf{x}(\sigma_{i})$ to $\mathbf{x}(\sigma_{i+1})$ is:
  1. 1.
    Predict denoised data: $\hat{\mathbf{x}}_{0}=F_{\theta}(\mathbf{x}(\sigma_{i});\sigma_{i})$.
  2. 2.
    Calculate drift: $\mathbf{d}=(\mathbf{x}(\sigma_{i})-\hat{\mathbf{x}}_{0})/\sigma_{i}$.
  3. 3.
    Take step: $\mathbf{x}(\sigma_{i+1})=\mathbf{x}(\sigma_{i})+\mathbf{d}(\sigma_{i+1}-\sigma_{i})$.
More sophisticated samplers are often used to improve quality and efficiency [41, 73]. Several applications require guiding generation with a condition $\mathbf{y}$ to sample from $p_{\theta}(\mathbf{x}|\mathbf{y})$. This is often done by feeding an embedding of $\mathbf{y}$ into the network $D_{\theta}(\mathbf{x}(\sigma),\sigma,\mathbf{y})$. However, it requires the conditions to be known a priori for training.
2.4 Diffusion Posterior Sampling and Data Assimilation
Another method to sample from the posterior $p_{\theta}(\mathbf{x}|\mathbf{y})$ is by combining a pre-trained generative (unconditional) prior $p_{\theta}(\mathbf{x})$ with a measurement likelihood $p(\mathbf{y}|\mathbf{x})$. From Bayes’ theorem, the posterior score is the sum of the prior and likelihood scores:
$\nabla_{\mathbf{x}_{t}}\log p_{\theta}(\mathbf{x}_{t}|\mathbf{y})\approx\nabla_{\mathbf{x}_{t}}\log p_{\theta}(\mathbf{x}_{t})+\nabla_{\mathbf{x}_{t}}\log p(\mathbf{y}|\mathbf{x}_{t})$ (8)
This formulation is structurally analogous to guidance methods like CFG [74], where the prior score (from the unconditional model) is modified by a guidance term (the difference between conditional and unconditional scores). Here, the guidance term is the gradient of the log-likelihood. A key challenge is estimating this likelihood score $\nabla_{\mathbf{x}_{t}}\log p(\mathbf{y}|\mathbf{x}_{t})$.
For a differentiable measurement model $M$ and Gaussian observation noise $\Sigma_{y}$, DPS [46] approximates the likelihood $p(\mathbf{y}|\mathbf{x}_{t})$ by first estimating the clean data $\hat{\mathbf{x}}_{0}(\mathbf{x}_{t},t)$ from the noisy sample $\mathbf{x}_{t}$ and then evaluating the likelihood at this estimate:
$p(\mathbf{y}|\mathbf{x}_{t})\approx\mathcal{N}(\mathbf{y}|M(\hat{\mathbf{x}}_{0}(\mathbf{x}_{t},t)),\Sigma_{y})$ (9)
The estimate $\hat{\mathbf{x}}_{0}$ is obtained via Tweedie’s formula [75], which connects the posterior mean to the score of the noisy data distribution:
$\hat{\mathbf{x}}_{0}(\mathbf{x}_{t},t)=\mathbb{E}[\mathbf{x}_{0}|\mathbf{x}_{t}]\approx\frac{\mathbf{x}_{t}+\sigma_{t}^{2}\mathbf{s}_{\phi}(\mathbf{x}_{t},t)}{\mu_{t}}$ (10)
where $\mathbf{s}_{\phi}$ is the learned score model, and $\mu_{t},\sigma_{t}$ are scaling factors. Score-based Data Assimilation (SDA) [76] refines this by accounting for the variance of the estimate $\hat{\mathbf{x}}_{0}$, yielding a more stable perturbed likelihood:
$p(\mathbf{y}|\mathbf{x}_{t})\approx\mathcal{N}\left(\mathbf{y}\mid M(\hat{\mathbf{x}}_{0}),\Sigma_{y}+\frac{\sigma_{t}^{2}}{(\mu_{t})^{2}}\mathbf{J}_{M}(\hat{\mathbf{x}}_{0})\mathbf{\Gamma}\mathbf{J}_{M}(\hat{\mathbf{x}}_{0})^{T}\right)$ (11)
where $\mathbf{J}_{M}$ is the Jacobian of the measurement function and $\mathbf{\Gamma}$ is the term depends on the eigen-decomposition of $\Sigma_{x}$. In practice, to simplify the approximation, the term $\mathbf{J}_{M}\mathbf{\Gamma}\mathbf{J}_{M}^{T}$ is replaced by a constant (diagonal) matrix (refer section 3.2 in [76]).
2.5 Diffusion Data Assimilation for EDM
In our work, we adapt the data assimilation process for the EDM framework. As the network $F_{\theta}$ is trained to predict $\mathbf{x}_{0}$, the score is given by Eq. 4. In the SDE convention used by EDM ($\mu_{t}=1$), substituting this into Tweedie’s formula (Eq. 10) gives a direct and intuitive estimate for the clean data:
$\hat{\mathbf{x}}_{0}=F_{\theta}(\mathbf{x}_{t};\sigma_{t})$ (12)
The likelihood score is then computed based on the SDA approximation (Eq. 11), simplified for the EDM case:
$\mathbf{s}_{l}(\mathbf{x}_{t},\sigma_{t};\mathbf{y})=\nabla_{\mathbf{x}_{t}}\log p(\mathbf{y}|\mathbf{x}_{t})\propto\nabla_{\mathbf{x}_{t}}\frac{-||\mathbf{y}-M(F_{\theta}(\mathbf{x}_{t};\sigma_{t}))||^{2}}{\Sigma+\hat{\Gamma}\sigma_{t}^{2}}$ (13)
where $\Sigma$ is the variance of noise observation and $\hat{\Gamma}$ approximates $\mathbf{J}_{M}\mathbf{\Gamma}\mathbf{J}_{M}^{T}$. These hyperparameters are constants that are tuned based on the precision of measurements and desired accuracy. This likelihood gradient is then used to guide the standard EDM sampling process. The first-order update step becomes:
  1. 1.
    Predict clean data: $\hat{\mathbf{x}}_{0}=F_{\theta}(\mathbf{x}_{\sigma_{i}};\sigma_{i})$.
  2. 2.
    Calculate prior drift: $\mathbf{d}_{\text{prior}}=(\mathbf{x}_{\sigma_{i}}-\hat{\mathbf{x}}_{0})/\sigma_{i}$.
  3. 3.
    Evaluate likelihood gradient: $\mathbf{s}_{l}(\mathbf{x}_{\sigma_{i}},\sigma_{i};\mathbf{y})$ using Eq. 13.
  4. 4.
    Calculate posterior drift: $\mathbf{d}_{\text{pos}}=\mathbf{d}_{\text{prior}}-\lambda_{g}\cdot\sigma_{i}\cdot\mathbf{s}_{l}$, where $\lambda_{g}$ is a guidance scale and $\sigma_{i}$ is multiplied to correctly incorporate it into the drift term of the probability flow ODE.
  5. 5.
    Perform sampling step: $\mathbf{x}_{\sigma_{i+1}}=\mathbf{x}_{\sigma_{i}}+\mathbf{d}_{\text{pos}}\cdot(\sigma_{i+1}-\sigma_{i})$.
This iterative procedure generates samples that adhere to both the learned data prior and the provided observations $\mathbf{y}$. Finally, we modify the stochastic sampler proposed in the EDM paper [41] to define Algorithm 1, which we employ in this work.
Algorithm 1 Stochastic Posterior Sampler
1:Require: Pre-trained denoiser $D_{\theta}(\mathbf{x};t)$, time steps $\{t_{i}\}_{i=0}^{N}$ where $t_{N}=0$, observation $\mathbf{y}$, guidance scale $\lambda_{g}$.
2:Require: Churn schedule parameters $S_{\text{churn}},S_{t_{\min}},S_{t_{\max}}$.
3:Sample $\mathbf{x}_{t_{0}}\sim\mathcal{N}(0,t_{0}^{2}\mathbf{I})$
4:for $i=0,\dots,N-1$ do
5:     $\gamma_{i}\leftarrow\min\left(\frac{S_{\text{churn}}}{N},\sqrt{2}-1\right)\text{ if }t_{i}\in[S_{t_{\min}},S_{t_{\max}}]\text{ else }0$triangleright Determine amount of noise injection (churn)
6:     $\hat{t}_{i}\leftarrow t_{i}+\gamma_{i}t_{i}$
7:     Sample $\boldsymbol{\epsilon}_{i}\sim\mathcal{N}(0,\mathbf{I})$
8:     $\mathbf{x}_{\hat{t}_{i}}\leftarrow\mathbf{x}_{t_{i}}+\sqrt{\hat{t}_{i}^{2}-t_{i}^{2}}\boldsymbol{\epsilon}_{i}$triangleright Add noise to move from $t_{i}$ to $\hat{t}_{i}$
9:     $\hat{\mathbf{x}}_{0}^{\text{pred}}\leftarrow D_{\theta}(\mathbf{x}_{\hat{t}_{i}};\hat{t}_{i})$
10:     $\mathbf{d}_{\text{prior}}^{\text{pred}}\leftarrow(\mathbf{x}_{\hat{t}_{i}}-\hat{\mathbf{x}}_{0}^{\text{pred}})/\hat{t}_{i}$
11:     $\mathbf{s}_{l}^{\text{pred}}\leftarrow\nabla_{\mathbf{x}_{\hat{t}_{i}}}\log p(\mathbf{y}|\mathbf{x}_{\hat{t}_{i}})$triangleright Evaluate likelihood gradient using Equation 13
12:     $\mathbf{d}_{\text{pos}}^{\text{pred}}\leftarrow\mathbf{d}_{\text{prior}}^{\text{pred}}-\lambda_{g}\cdot\hat{t}_{i}\cdot\mathbf{s}_{l}^{\text{pred}}$
13:     $\mathbf{x}_{t_{i+1}}^{\text{pred}}\leftarrow\mathbf{x}_{\hat{t}_{i}}+\mathbf{d}_{\text{pos}}^{\text{pred}}\cdot(t_{i+1}-\hat{t}_{i})$triangleright Initial Euler step
14:     if $t_{i+1}\neq 0$ then
15:         $\hat{\mathbf{x}}_{0}^{\text{corr}}\leftarrow D_{\theta}(\mathbf{x}_{t_{i+1}}^{\text{pred}};t_{i+1})$
16:         $\mathbf{d}_{\text{prior}}^{\text{corr}}\leftarrow(\mathbf{x}_{t_{i+1}}^{\text{pred}}-\hat{\mathbf{x}}_{0}^{\text{corr}})/t_{i+1}$
17:         $\mathbf{s}_{l}^{\text{corr}}\leftarrow\nabla_{\mathbf{x}_{t_{i+1}}^{\text{pred}}}\log p(\mathbf{y}|\mathbf{x}_{t_{i+1}}^{\text{pred}})$triangleright Re-evaluate likelihood gradient using Equation 13 (optional)
18:         $\mathbf{d}_{\text{pos}}^{\text{corr}}\leftarrow\mathbf{d}_{\text{prior}}^{\text{corr}}-\lambda_{g}\cdot t_{i+1}\cdot\mathbf{s}_{l}^{\text{corr}}$
19:         $\mathbf{x}_{t_{i+1}}\leftarrow\mathbf{x}_{\hat{t}_{i}}+\frac{1}{2}(\mathbf{d}_{\text{pos}}^{\text{pred}}+\mathbf{d}_{\text{pos}}^{\text{corr}})\cdot(t_{i+1}-\hat{t}_{i})$triangleright 2nd-order correction step
20:     else
21:         $\mathbf{x}_{t_{i+1}}\leftarrow\mathbf{x}_{t_{i+1}}^{\text{pred}}$
22:     end if
23:end for
24:return $\mathbf{x}_{t_{N}}$
3 Experimental configuration
3.1 Motivation
Data assimilation for atmospheric science has traditionally been performed by using a short-term forecast from a numerical-based model (e.g., weather forecasting or climate models) as the prior, which is then updated with observations using numerical techniques (e.g., Kalman filtering or variational methods) to provide a best estimate of the atmospheric (or ocean, land surface, etc.) state at a particular time on a regular grid [77]. This helps overcome the problem with observations (e.g., from satellites, weather stations, buoys) that are irregular in time and space, noisy, and often incomplete. Data assimilation combines the strengths of both: observational accuracy with model-based physical consistency and allows for the estimation of the state on a regular grid. This approach has been successful and sits at the heart of modern-day weather prediction. One particularly important use case of data assimilation has been the ability to generate physically consistent, comprehensive, and temporally continuous depiction of the Earth system, call reanalysis (e.g., ERA5) [23]. However, classical methods for generating reanalysis data from observations and a numerical model tend to be computationally expensive require significant use of high-performance computing resources. This can be seen with ERA5 which required massive amounts of computational resources over multiple years to produce the nearly 80 years of reanalysis data [23]. If machine learning could produce similar quality estimates of the atmosphere using observations, this could represent orders of magnitude improvement in computational efficiency for data assimilation.
3.2 Data
We use ERA5, a state-of-the-art global reanalysis dataset produced by the European Centre for Medium-Range Weather Forecasts (ECMWF), which provides hourly estimates of a large number of atmospheric, land, and oceanic climate variables at a resolution of  31 km [23]. For this work, we use the 1.40525° (128 × 256 grid points) ERA5 dataset from WeatherBench [78], which used bilinear interpolation to regrid from the native resolution to 1.4°. The choice to use a lower resolution version of ERA5 was due to the computational limitations of training an unconditional diffusion model without large scale compute access.
For observational data, we rely on the Integrated Global Radiosonde Archive (IGRA), a comprehensive dataset of weather balloon observations from several stations worldwide, providing upper-air measurements like temperature, humidity, geopotential, pressure, and wind at different pressure levels [79]. Furthermore, we also construct an observation data set that is a further coarsened ERA5 at 11.242°(16x32 grid points) using bicubic interpolation. This is mentioned as LR further in the results.
3.3 Measurement Operators
For our super-resolution tasks, the measurement operation involves a downsampling operator that reduces the resolution of the data using interpolation techniques such as bicubic filtering. These downsampled values are compared with the available low-resolution data and used to guide our diffusion model. We mention these samples as Sample-LR. In contrast, our data assimilation tasks involve irregular or sparse observational locations. Our IGRA operator interpolates values at arbitrary spatiotemporal coordinates (e.g., latitude and longitude) using grid-based sampling. It defines a non-uniform measurement operator $M$ that evaluates the high-resolution field at the queried coordinates via bicubic interpolation in normalized coordinate space. This enables the comparison between model predictions and sparse ground truth measurements, forming the basis of our diffusion data assimilation frameworks. We mention these samples as Sample-IGRA. For our multimodal data assimilation we compute the misfit for both low-resolution and the IGRA measurements and use a weighted combination of the misfits to guide our diffusion sampling. We mention these samples as Sample IGRA+LR henceforth. For instance, the operator $M$ in Equation 13 can be the IGRA operator, which takes the generated full field as input and returns the corresponding values at the locations where we have the true IGRA measurements $\mathbf{y}$.
(a) 2m Temperature (K)
(b) U-Wind (m/s)
(c) V-Wind (m/s)
(d) Specific humidity (kg/kg)
Figure 1: Super-resolution using zero-shot posterior sampling given structured grid low-resolution observations (sample-LR). Contours showing mean of samples at specific time instances.
3.4 Network architecture and hyperparameters
[…]
4 Results and discussion
We outline results from our various experiments in the following. The training of the score-based generative model required approximately 30 hours on 16 Nvidia A100 GPUs, while generating a sample conditionally or with observation of sparse information requires only 10 seconds on 1 A100 GPU.
4.1 Super-resolution from structured coarse-grained fields
Figure 2: Ensemble based mean (left) and standard deviations (right) obtained from ensembles obtained with sample-LR zero-shot sampling. Showing 2m temperature, 10m u component of wind, 10m v component of wind, and specific humidity at 1000mb pressure from top to bottom.
(a) 2m Temperature
(b) 10m u-wind
(c) 10m v-wind
(d) 1000mb specific humidity
Figure 3: Spectral recovery exhibited by zero-shot sampling from ERA5 generative model when using sample-LR for zero-shot sampling.
Figure 4: Traces for reconstructed flow-fields (including mean and standard deviation) compared with ground truth using sample-LR for zero-shot sampling.
In the first experiment, we assess the ability to reconstruct the base-resolution ERA5 state given low-resolution observations from the same dataset (sample-LR). We qualitatively assess the reconstruction ability via test-time score-matching using visualizations of the fully recovered flow-fields for some selected variables in Figure 1 where coarse-grained sparse observations as well as the ground truth flow-fields at a particular instant are available. Error contours are obtained from using the ensemble mean as the predicted full state recovery. The contours indicate the ability to recover fine-scaled features using the score-matching technique effectively. We also assess the mean and variance of the generated samples from the score-matching procedure in Figure 2 where we can see correlations between variables and regions of high or low uncertainty. In particular, for recovering the 2m-temperature variable, we observe a higher uncertainty over the continents in contrast with wind speeds at 10m heights. For recovering specific humidity at 1000mb, we observe no coherent regions of high uncertainty aside from the general region of the lower latitudes where high values of specific humidity are commonly observed.
Quantitative analyses are also performed by looking at the quality of spectral recovery from angle-averaged kinetic energy spectra for the various variables as shown in Figure 3. In these plots, it can be seen that the recovered spectra from the zero-shot score-matching procedure leads to good agreement with the ground truth spectra for different variables all the way through to the cut-off wavenumber. Finally, we also assess the spatial variability of the generated fields by looking at traces along select latitudes across longitude in Figure 4. Here, the ensemble mean and variance are also plotted showing certain regions with higher uncertainty. In general, a good recovery of the trends is observed, including the capture of sharp peaks and troughs accurately across different variables.
4.2 Super-resolution from unstructured data and different modality
(a) Temperature (K)
(b) U-wind (m/s)
(c) V-wind (m/s)
Figure 5: Super-resolution using zero-shot sampling with posterior updates based on IGRA observations (sample-IGRA). The rows indicate ensemble mean (top), ground truth (middle), and ensemble standard deviation from multiple zero-shot samples. Continents show lower uncertainty since IGRA radiosondes are predominantly based on land.
Our next experiment is to extend the setting of ERA5 state recovery given sparse and unstructured observations from a source of data that is different to that used for training. This can be contrasted to the previous section, where it was assumed that sparse observations were obtained from the original training data set distribution. Specifically, we construct a problem statement where our unconditional ERA5 generation is informed, during test, using sparse observations of the IGRA radiosonde dataset (i.e., the sample-IGRA configuration). Notably, this dataset is available only at land locations and therefore represents a sensing platform that is constrained. Figure 5 shows sparse locations where IGRA readings are sampled and their corresponding full-state recovery performance for exemplar time instances. Clear patterns may be distinguished for uncertainty estimates. In particular, one can observed much lower uncertainty over the land which corresponds to regions where sensor measurements are available in real-time. Higher uncertainty, particularly for wind-speeds are observed over the oceans.
(a) Temperature
(b) U-Wind
(c) V-Wind
(d) Temperature
(e) U-Wind
(f) V-Wind
Figure 6: Scatters at IGRA sensor locations (top) and non-IGRA locations (bottom) from zero-shot sampling using sample-IGRA. The knowledge of IGRA sensor measurements improves the reconstruction at sensor locations significantly as seen above. For locations where observations are not available, some variables, such as temperature are reconstructed more accurately than others such as wind speeds.
(a) Geopotential 200
(b) Specific humidity 1000
(c) Temperature 150
Figure 7: Scatters for additional fields at non-IGRA locations using zero-shot sampling from sample-IGRA. The knowledge of IGRA sensor measurements improves the reconstruction at sensor locations significantly as seen above.
(a) 2m Temperature
(b) 10m u-wind
(c) 10m v-wind
(d) 1000mb specific humidity
Figure 8: Spectral recovery exhibited by zero-shot sampling from ERA5 generative model when observing sparse observation data from the IGRA dataset (sample-IGRA). Slight cut-off wavenumber errors are observed for the reconstructed flow-fields.
Figure 9: Traces for reconstructed flow-fields (including mean and standard deviation) compared with ground truth for the sample-IGRA zero-shot sampling.
In Figure 6, we plot scatters for three representative variables at IGRA locations, as generated by the score-matching procedure. We expect to see the values generate collapse on a 45 line indicating that IGRA observations have been assimilated well. We also show the ground truth ERA5 values at the specific day in these scatters, where significant deviations are observed for the wind speeds. This indicates that the state recovery procedure has prioritized the likelihood obtained by IGRA observations at this point. In the same figure, we assess the performance of the generative model (after score-matching) for recovering the ground truth ERA5 at non-sampled locations given IGRA observations. Here it is seen that the spread for the generated values at non-IGRA points is generally reduced (in particular for temperature and specific humidity), indicating the IGRA observations have helped constrain the generator to the specific state. The spread from the unconditional samples indicates that while the unconditional generative process is able to provide fields that are ERA5-like, they do not recover the specific flow-fields without score-matching.
As in the previous section, we also assess the spectral quality of the ERA5 state recovery, given IGRA observations, through angle-averaged kinetic energy spectra as shown in Figure 8. In general, similar recovery quality is observed with a slight increase in grid cut-off wavenumber errors for the ensemble means. We also draw the reader’s attention to the higher spread in the variance for wind-speed reconstruction indicating how the sensing of IGRA observations may not necessarily correlate with high quality state or spectral recovery for all variables. We believe this to be a promising metric to optimize for improved sensing platforms. Finally, we also plot traces of the generated state recovery for different latitudes across longitudes in Figure 9, where comparisons with the ground truth ERA5 are also made. One can observe an increased variance from the samples of the generative model indicating greater uncertainty in the state reconstruction. This reflects the intrinsic challenges of reconstructing the reanalysis from partial and spatially anisotropic observations such as IGRA.
4.3 Multimodal super-resolution
(a) 2m Temperature
(b) 10m u-wind
(c) 10m v-wind
(d) 1000mb specific humidity
Figure 10: Spectral recovery exhibited by zero-shot sampling from ERA5 generative model when observing sparse observation data from the IGRA dataset as well as from low-resolution reanalysis (sample IGRA+LR). The biases in the higher wavenumbers obtained from state reconstructions with only IGRA observations are reduced by using low-resolution observations as well.
(a) Surface Temperature
(b) U-Wind
(c) V-Wind
Figure 11: Scatter, at IGRA sensing locations, from an exemplar reconstruction using zero-shot score matching when using the sample IGRA+LR configuration. At these locations, one can observe a closer agreement to the IGRA values.
Figure 12: Traces for zero-shot reconstructed flow-fields (including standard deviation as shaded regions) compared with ground truth for state reconstruction with sample-LR, sample-IGRA, and sample-IGRA+LR configurations. The results indicate that the addition of low-resolution ERA5 observations improve those from solely IGRA based reconstructions.
(a) Surface Temperature
(b) U-Wind
(c) V-Wind
Figure 13: Scatter, at non-IGRA locations, from an exemplar reconstruction using zero-shot score matching when using the sample IGRA+LR configuration. At these locations, one can observe a closer agreement to the ERA5 values indicate effective multimodal data assimilation.
In this section, we evaluate one of the greatest strengths of zero-shot sampling from generative models, namely, their ability to assimilate multiple sources of data during test time. Specifically, we perform experiments that rely on observations of both the low-resolution ERA5 observations on a coarse and structured grid, as well as IGRA observations on the unstructured data (previously introduced as the sample-IGRA+LR configuration). In Figure 10 we show the spectral recovery properties of the multifidelity assimilation versus those obtained solely from IGRA and the structured grid observations. One of the key benefits of using data from the low-resolution ERA5 in addition to IGRA is the significant reduction in uncertainty in the lower wavenumbers for the wind-speed spectra in addition to a correction of the high-wavenumber errors obtained when solely using IGRA. This is also observed in longitudinal traces in Figure 12 where the variance in samples is reduced given more relevant information. When assessing the scatter of the generated values for exemplar state recovery at the IGRA sensor locations, as shown in Figure 11, we observe that the generative model is able to sample in a manner that is closer to the observed IGRA magnitudes despite using both low-resolution ERA5 and IGRA for score-matching. This indicates an ability to balance the influence of various sources of data based on the proximity to their observation locations. This is reinforced by evidence in Figure 7, where scatter plots for reconstructions at non-IGRA sensing locations indicate an improved reconstruction of ERA5 now that coarse-grid ERA5 observations are provided during zero-shot score matching. We remark that while using low-resolution observations of reanalysis represents an unrealistic source of data during real-time data fusion, it represents a proof-of-concept of fusing another high-quality source of data to reduce uncertainties during assimilation.
(a) 2m Temperature (K)
(b) U-Wind (m/s)
(c) V-Wind (m/s)
(d) Specific humidity (kg/kg)
Figure 14: An assessment of the difference between zero-shot score-based sampling and the ground truth for the various sampling configurations. We observe that the sample-IGRA performance, while competitive for temperature and specific humidity reconstructions, need additional observations from coarse-grid ERA5 for improved wind-speed reconstructions. This provides evidence that combinations can be used for improved reconstructions in zero-shot flow matching.
5 Discussion and Conclusion
Rapid assimilation of observations from real-time platforms has the potential to provide significant gains in the accurate assessment of high-dimensional dynamical systems. The motivation of this study was to propose avenues for reducing the significant computational costs for achieving the same using classical numerical methods for data assimilation of high-dimensional multiscale systems. To that end, this study proposed the construction of an unconditional generative model that is trained to match the score of the underlying probability distribution from which arbitrary training data may be sampled. Once this score is approximated with a neural network, a zero-shot score-matching technique built on a Bayesian formulation can be used to rapidly assimilate sparse multifidelity observational data with negligible real-time costs. In this study, we construct an experiment to recover the full state of a reanalysis dataset given partial observations from a coarse-grid observation of the same as well as when sparse and spatially unstructured data are obtained from a different dataset altogether. In both such examples, we observe that the score-matching technique is proficient in recovering the original state of the reanalysis data on the fine grid. Notably, when both sets of partial observations are used for a multifidelity state recovery exercise, we observe that the data fusion strategy is able to balance proximity to both data sets based on their spatial sensing locations. This indicates that zero-shot score matching can be used as an effective and computationally inexpensive tool to accurately perform data fusion from data of varying sources and spatiotemporal fidelity. Based on the promising results from this study, future extensions of this framework include the integration of dynamical considerations, i.e., performing state vector trajectory reconstructions given time-series traces at observation locations, integrating both Eulerian and Lagragian observations within a common fusion paradigm, and performing optimal sensor placement for high quality state recovery.
Acknowledgments
This research used resources of the Argonne Leadership Computing Facility, which is a U.S. Department of Energy Office of Science User Facility operated under contract DE-AC02-06CH11357. DC, HG, and RM acknowledge support from DOE ASCR award “Inertial neural surrogates for stable dynamical prediction" and DOE FES award “DeepFusion Accelerator for Fusion Energy Sciences in Disruption Mitigation". DC, HG, RM also acknowledge the support of computing resources through an AI For Science allocation at the National Energy Research Scientific Computing Center (NERSC) and from Penn State Institute for Computational and Data Sciences.